# Long-Range Interacting Particles on a Helix: A Statistical and Correlation Analysis of Equilibrium Configurations

### v1

@@ I Introduction
Driven by the functionalities of helices in nature, fabrication methods and properties of helical nanostructures have been studied in detail [12]. In addition to fully guided manufacturing methods [13], the research has yielded self-assembled helical structures, for instance, due to crystal growth [13, 14, 15, 12, 16], template-assisted approaches [13, 17], or long-range ordering on cylindrical surfaces [18, 19, 20]. Many of these nanohelices possess remarkable mechanical, electrical, and optical properties [21, 14, 22, 23]. Notable examples include the superelasticity of ZnO nanohelices [21] or the unique capacity of coiled carbon nanotubes for electronic property modification based on chirality, which may offer potential applications in nanoelectronic devices [14]. An additional, more fundamental approach has been realized in the form of a double-helical trap for cold neutral atoms. This trap utilizes an optical nanofiber surrounded by a two-color evanescent field, thereby establishing a platform for fundamental physical research [24].
@@ II Setup, Hamiltonian and two-body problem
where $R$ is the radius, $h$ is the pitch of the helix, and $\phi_{i}$ are its parametric coordinates (azimuthal angles). In Cartesian coordinates, the Hamiltonian has the form $\mathcal{H}=\mathcal{T}+\mathcal{V}$, where $\mathcal{T}=\sum_{i}\frac{m_{i}}{2}\dot{\mathbf{r}}_{i}^{2}$ is the kinetic energy and $\mathcal{V}=\sum_{i<k}g_{ik}/\left|\mathbf{r}_{i}-\mathbf{r}_{k}\right|$ is the potential energy. We consider equal masses $m_{i}=m$ and equal coupling constants $g_{ik}=g=q^{2}/(4\pi\varepsilon_{0})$. The arc length parametrization
$s_{i}:\phi_{i}\mapsto\int_{0}^{\phi_{i}}\left|\partial\mathbf{r}_{i}(\phi_{i}^{\prime})/\partial\phi_{i}^{\prime}\right|\,\mathrm{d}\phi_{i}^{\prime}$ (2)
@@
Rather than employing path lenghts, the more intuitive quantitative measure, $\Phi_{i}=\tilde{s}_{i}/(2\pi\xi)$, which counts the number of helical windings traversed, will be utilized. For the sake of simplicity, we also utilize the dimensionless potential energy $V=\mathcal{V}R/g$.
Figure 2: (a) Potential energy $V$ for two charged particles confined on a helix, depending on the interparticle distance $\Phi=\Phi_{1}$. The potential is shown for $\rho=0.2$, $\rho=0.5$, $\rho=0.8$, and $\rho=1.1$. Additionally, the dashed lines show the upper and lower envelope of the potential energy for $\rho=0.2$ and $\rho=1.1$, respectively. (b) An intersection of the many-body potential landscape $V(\Phi)$, where $\Phi=\Phi_{1}=\Phi_{2}=\ldots=\Phi_{N-1}$, i.e. for equidistant particles, for different $N$ and $\rho=0.8$. The energetically lowest curve corresponds to $N=2$, the second lowest to $N=3$, etc. The black crosses mark the minimum of the outermost potential well of each curve.
@@
In the exact limit of $\rho=h=0$, the helix becomes a circle, resulting in ECs on a polygon with equal interparticle distances. For a helix with finite $h$ the ECs are approximately odd multiples of half a helical winding i.e. we have $\Phi=n+\frac{1}{2}$, $n\in\mathbb{N}_{0}$. In these ECs, the particles are located approximately on opposite sides of the windings thereby maximizing the Euclidean distance and consequently minimizing the potential energy. To be more precise, the true minima of the potential energy are located at values slightly greater than odd multiples of half a helical winding due to the helical winding along its axis. Fig. 2(a) illustrates the oscillatory behavior of the potential energy for the case of two particles $N=2$ for different values of the parameter $\rho$, with each minimum corresponding to an EC. The number of ECs scales approximately as $\lfloor 2\pi/\rho^{2}\rfloor$, i.e. a larger ratio $\rho$ leads to a larger number of stable ECs. For $\Phi>2\pi/\rho^{2}$, the quadratic term in the denominator of Eq. (4) becomes predominant. Consequently, there are no further minima leading to stable ECs. For the ECs near that cutoff value, the deviation from the aforementioned distances of approximately odd multiples of half a helical winding is significant, thereby establishing a kind of spatial irregularity for the distances between the ECs. By increasing $\rho$, this irregularity becomes dominant.
@@
The investigation of the equilibrium structures of helical few-body systems in the range $N=2-8$, conducted in this work, will further manifest their hierarchical properties between regularity and irregularity with a focus on their statistical behaviour. To accomplish this we choose for $\rho$ the value $0.8$. Let us provide a teaser of what awaits us. Fig. 2(b) shows an one-dimensional intersection of the $(N-1)$-dimensional potential landscape for equidistant particles on the helix, i.e. along its global diagonal $\Phi=\Phi_{1}=\Phi_{2}=\ldots=\Phi_{N-1}$ and for different numbers of particles. Despite the fact that this illustration is limited to a low-dimensional intersection of the originally high-dimensional potential energy surface and does not encompass any ECs of the corresponding few-body systems, it offers a valuable insight into the structure of the potential landscape. It is evident that the addition of particles to the system results in an increase not only of the mere potential energy but in particular also of the complexity of the potential energy, attributable to the cumulative Coulomb interaction of the particles. Of particular significance is the observation that the overall structure extends on what we have already been finding for the two-body system: it exhibits (i) an overall oscillatory potential (ii) a hierarchical organization of energies and (iii) spatial irregularities in the minima/wells of the intersection potential.
@@
Tentatively, the wells of the intersections (see Fig. 2(b)) for small values of $\Phi$ show a highly irregular oscillatory behaviour. It is noteworthy that in general the potential wells become deeper upon transitioning to higher values of $N$. There exist however manifest exceptions. In Fig. 2(b), this is at least visible for the minima at $\Phi\approx 5/2$ and $\Phi\approx 7/2$. Increasing the number of particles for a given value of $\Phi$ (i.e. moving from the curve for $N$ to the curve for $N+1$ in Fig. 2(b)) results in a greater total extension $(\sum_{i=1}^{N-1}\Phi_{i})$ of the particle chain. Accordingly, the axial Coulomb forces exerted on the outermost particles (the first and the last particle of the particle chain) prevail for $N+1$ particles even at lower values of $\Phi$ compared to the $N$ particle chain. This ultimately leads to the flattening of the last potential well when $N$ is increased, as marked by the black crosses in Fig. 2(b). Nonetheless, this attribute is not entirely generic, as in the transition from $N=2$ to $N=3$, two minima disappear, and for $6\leq N\leq 8$, respectively $9\leq N\leq 14$, the last minimum is stable. It is imperative to reiterate that Fig. 2(b) is merely an intersection and does not embody the high-dimensional potential landscape. Nevertheless, this intersection partially reflects the structural peculiarities of the potential landscape and its local equilibria, as we shall see in the subsequent sections.
@@ III Methodology and computational approach
The statistical and correlation analysis of the spatial and energetic properties of the ECs of a long-range interacting particle system on a helix necessitates the finding of all local minima of our high-dimensional oscillatory potential energy surface (see Eq. (4)). This is illustrated in Fig. 2(b) and will be further elaborated below: the potential energy landscape shows a mix of regular and irregular oscillations and undergoes substantial fluctuations across various particle configurations while maintaining a hierarchical structure. The exponential growth in configuration space volume, concomitant with the rise in dimensionality, poses a major challenge in identifying all local equilibria of the potential landscape. We address this by implementing a multi-start approach, entailing the definition of multiple sampling points. Each sampling point is then utilized as an initial condition for a local optimizer. Notwithstanding the aforementioned ”disorder” inherent in the potential landscape, the repulsion within and across helical windings must engender approximate regularity in the occurrence of ECs, which motivates the definition of the sampling points as a periodic hyperlattice $L\subseteq Q$ in the configuration space $Q\subseteq\mathbb{R}^{N-1}$. Every point in this space is given by a vector $\mathbf{\Phi}=(\Phi_{1},\Phi_{2},\ldots,\Phi_{N-1})$ of the internal coordinates. The hyperlattice is defined as $L=\tilde{\Phi}_{1}\times\tilde{\Phi}_{2}\times\ldots\times\tilde{\Phi}_{N-1}$, where the subsets $\tilde{\Phi}_{k}\subseteq\mathbb{R}$, with $k=1,2,\ldots,N-1$, define $n$ evenly spaced points in the $k$-th dimension in a given interval $[a,b]\subseteq\mathbb{R}$. The optimization algorithm employed is L-BFGS-B [36, 37], which implements a quasi-Newton method with bound constraints for the $k$-th coordinate.
@@
The computational effort resulting from this approach is increasing exponentially with the number of particles employed in the system ($\propto n^{N-1}$). Furthermore, the highly oscillatory structure of the potential landscape requires a high sampling point density $\delta=n/(b-a)$ to find all local equilibria. Reducing the search space is one way to lower the computing costs under these conditions. As indicated in the previous sections, the maximum value of $\Phi_{k}$ that allows for ECs either remains approximately constant or decreases when a particle is added to the chain. Consequently, this maximum value serves as an upper bound constraint $b$ for the system with an incrementally larger number of particles. On the other hand, the inversion of the helix with a given configuration $\mathbf{\Phi}$ results in an identical configuration $-\mathbf{\Phi}$ with the same potential energy. Therefore, the search space can be further reduced by approximately half through the elimination of these identical configurations. We accomplish this by employing a recursive constraint for $\Phi_{i}$ and $\Phi_{N-i}$, $i=1,\ldots,k=\lfloor(N-1)/2\rfloor$ and $N\geq 3$. We enforce that $\Phi_{1}\geq\Phi_{N-1}$; $\Phi_{2}\geq\Phi_{N-2}$ if $\Phi_{1}=\Phi_{N-1}$; $\Phi_{3}\geq\Phi_{N-3}$ if $\Phi_{1}=\Phi_{N-1}\land\Phi_{2}=\Phi_{N-2}$; $\ldots$ ; $\Phi_{k}\geq\Phi_{N-k}$ if $\Phi_{1}=\Phi_{N-1}\land\,\ldots\,\land\Phi_{k-1}=\Phi_{N-(k-1)}$. Formally, this constraint can be written as a lexicographical ordering of the subsets $M=\left\{\Phi_{1},\ldots,\Phi_{k}\right\}$ and $N=\left\{\Phi_{N-1},\ldots,\Phi_{N-k}\right\}$: $M\geq_{\mathrm{lex}}N$. In our optimization algorithm, points of $L$ that invalidate this constraint are introduced into the interstitial voids of the lattice structure.
@@ IV.1 Spatial equilibrium configurations: Statistics and correlations
Figure 3: The total number of attained ECs, contingent upon the number of particles. The inset illustrates the number of ECs by $N$ on a double logarithmic scale.
The analysis of the statistics and correlations of the spatial properties of the ECs will be conducted for each particle number in the range 2-8. In consideration of the denotation of an EC as $\mathbf{\Phi}_{0}=(\Phi_{0}^{1},\Phi_{0}^{2},\ldots,\Phi_{0}^{N-1})$, the distribution of the totality of the interparticle distances $\Phi_{0}^{i}$, $i=1,2,\ldots,N-1$ of all ECs will be examined, independent of their sequential ordering within an EC. Building upon this distribution, a mapping of these $\Phi_{0}^{i}$ to distinct clusters will be employed to elucidate the spatial structure of the ECs. This will be achieved by indexing the clusters and then correlating the $\Phi_{0}^{i}$ values within an EC based on their assigned cluster index. Finally, we will gain insight into the spatial structure by studying the distribution of ECs within the $(N-1)$-dimensional potential landscape, quantifying their Euclidean distances in the configuration space.
@@ IV.1.1 Distribution of interparticle distances
Figure 5: (a) The probability on how many distinct clusters the interparticle distances $\Phi_{0}^{i}$ of an EC (Clusters/EC) are distributed in dependence of the number of particles $N$. For example, for the EC $\mathbf{\Phi}_{0}\approx(1.35,0.36)$, $C(\Phi_{0}^{1})=2$ and $C(\Phi_{0}^{2})=1$, so $\Phi_{0}^{i}$ is distributed on two Clusters in that EC. The probability is normalized for each $N$. (b-h) The probability on which specific clusters the $\Phi_{0}^{i}$ are distributed, given the value of Clusters/EC. Note that $C(\Phi_{0}^{i})$ is cut at the value of 8 because $C(\Phi_{0}^{i})=9,10$ are only relevant for the trivial case of $N=2$. The probability is normalized for each value of Clusters/EC.
@@
A generally similar but more complex transition can be observed for the Clusters 1 and 2 (see Fig. 4(a-g)). While the maximum value of $\Phi_{0}^{i}$ in these custers also initially increases, the minimum value of $\Phi_{0}^{i}$ is decreasing, thereby forming a fine structure with two subclusters. For Cluster 2 the upper subcluster is broader than the lower one, evidenced by an asymmetry in the corresponding probability distributions. For $N>5$, cluster 2 decreases, predominantly by the fading of the lower subcluster. In contrast, cluster 1 exhibits a less pronounced dominance of the upper subcluster, characterized by a more equitable distribution of $\Phi_{0}^{i}$ across both clusters. This results in probability distributions, that are more or less symmetrical. Additionally, cluster 1 merely shows an increase in its width, while a subsequent decrease may only commence at higher values of $N$.

### v2

@@ I Introduction
Driven by the functionalities of helices in nature, fabrication methods and properties of helical nanostructures have been studied in detail [12]. In addition to fully guided manufacturing methods [13], the research has yielded self-assembled helical structures, for instance, due to crystal growth [13, 14, 15, 12, 16], template-assisted approaches [13, 17], or long-range ordering on cylindrical surfaces [18, 19, 20]. Many of these nanohelices possess remarkable mechanical, electrical, and optical properties [21, 14, 22, 23]. Notable examples include the superelasticity of ZnO nanohelices [21] or the unique capacity of coiled carbon nanotubes for electronic property modification based on chirality, which may offer potential applications in nanoelectronic devices [14]. An additional, more fundamental approach has been realized in the form of a double-helical trap for cold neutral atoms. This trap utilizes an optical nanofiber surrounded by a two-color evanescent field, thereby establishing a platform for fundamental physical research [24]. Furthermore, the arrangement of charged dust grains into quasi-crystal helical structures (zigzag, helix, interwoven helices) has been observed in a complex plasma in the context of external potentials and of self-confinement by a field of smaller background grains in a cylindrically symmetric setup [25, 26, 27].
@@ II Setup, Hamiltonian and two-body problem
where $R$ is the radius, $h$ is the pitch of the helix, and $\phi_{i}$ are its parametric coordinates (azimuthal angles). In Cartesian coordinates, the Hamiltonian has the form $\mathcal{H}=\mathcal{T}+\mathcal{V}$, where $\mathcal{T}=\sum_{i}\frac{m_{i}}{2}\dot{\mathbf{r}}_{i}^{2}$ is the kinetic energy and $\mathcal{V}=\sum_{i<k}g_{ik}/\left\lvert\mathbf{r}_{i}-\mathbf{r}_{k}\right\rvert$ is the potential energy. We consider equal masses $m_{i}=m$ and equal coupling constants $g_{ik}=g=q^{2}/(4\pi\varepsilon_{0})$. The arc length parametrization
$s_{i}:\phi_{i}\mapsto\int_{0}^{\phi_{i}}\left\lvert\partial\mathbf{r}_{i}(\phi_{i}^{\prime})/\partial\phi_{i}^{\prime}\right\rvert\,\mathrm{d}\phi_{i}^{\prime}$ (2)
@@
Rather than employing path lengths, the more intuitive quantitative measure, $\Phi_{i}=\tilde{s}_{i}/(2\pi\xi)$, which counts the number of helical windings traversed, will be utilized. For the sake of simplicity, we also utilize the dimensionless potential energy $V=\mathcal{V}R/g$.
Figure 2: (a) Potential energy $V$ for two charged particles confined on a helix, depending on the interparticle distance $\Phi=\Phi_{1}$. The potential is shown for $\rho=0.2$, $\rho=0.5$, $\rho=0.8$, and $\rho=1.1$. Additionally, the dashed lines show the upper and lower envelope of the potential energy for $\rho=0.2$ and $\rho=1.1$, respectively. (b) An intersection of the many-body potential landscape $V(\Phi)$, where $\Phi=\Phi_{1}=\Phi_{2}=\ldots=\Phi_{N-1}$, i.e. for equidistant particles, for different NN and $\rho=0.8$. The energetically lowest curve corresponds to $N=2$, the second lowest to $N=3$, etc. The black crosses mark the minimum of the outermost potential well of each curve.
@@
In the exact limit of $\rho=h=0$, the helix becomes a circle, resulting in ECs on a polygon with equal interparticle distances. For a helix with finite $h$ the ECs are approximately odd multiples of half a helical winding i.e. we have $\Phi=n+\frac{1}{2}$, $n\in\mathbb{N}_{0}$. In these ECs, the particles are located approximately on opposite sides of the windings thereby maximizing the Euclidean distance and consequently minimizing the potential energy. To be more precise, the true minima of the potential energy are located at values slightly greater than odd multiples of half a helical winding due to the helical winding along its axis. Fig. 2(a) illustrates the oscillatory behavior of the potential energy for the case of two particles $N=2$ for different values of the parameter $\rho$, with each minimum corresponding to an EC. The number of ECs scales approximately as $\lfloor 2\pi/\rho^{2}\rfloor$, i.e. a larger ratio $\rho$ leads to a larger number of stable ECs. Applying the condition $\mathrm{d}V/\mathrm{d}\Phi=0$ and $\lvert\sin(2\pi\Phi)\rvert\leq 1$, yields the upper limit $\Phi\leq 2\pi/\rho^{2}$, above which the quadratic term in the denominator of Eq. (4) becomes predominant. Consequently, there are no further minima leading to stable ECs. For the ECs near that cutoff value, the deviation from the aforementioned distances of approximately odd multiples of half a helical winding is significant, thereby establishing a kind of spatial irregularity for the distances between the ECs. By increasing $\rho$, this irregularity becomes dominant.
@@
The investigation of the equilibrium structures of helical few-body systems in the range N=28N=2-8, conducted in this work, will further manifest their hierarchical properties between regularity and irregularity with a focus on their statistical behaviour. To accomplish this we choose for $\rho$ the value $0.8$. Let us provide a teaser of what awaits us. Fig. 2(b) shows an one-dimensional intersection of the (N1)(N-1)-dimensional potential landscape for equidistant particles on the helix, i.e. along its global diagonal $\Phi=\Phi_{1}=\Phi_{2}=\ldots=\Phi_{N-1}$ and for different numbers of particles. Despite the fact that this illustration is limited to a low-dimensional intersection of the originally high-dimensional potential energy surface and does not encompass any ECs of the corresponding few-body systems, it offers a valuable insight into the structure of the potential landscape. It is evident that the addition of particles to the system results in an increase not only of the mere potential energy but in particular also of the complexity of the potential energy, attributable to the cumulative Coulomb interaction of the particles. Of particular significance is the observation that the overall structure extends on what we have already been finding for the two-body system: it exhibits (i) an overall oscillatory potential (ii) a hierarchical organization of energies and (iii) spatial irregularities in the minima/wells of the intersection potential.
@@
Tentatively, the wells of the intersections (see Fig. 2(b)) for small values of $\Phi$ show a highly irregular oscillatory behaviour. It is noteworthy that in general the potential wells become deeper upon transitioning to higher values of $N$. There exist however manifest exceptions. In Fig. 2(b), this is at least visible for the minima at $\Phi\approx 5/2$ and $\Phi\approx 7/2$. Increasing the number of particles for a given value of $\Phi$ (i.e. moving from the curve for $N$ to the curve for $N+1$ in Fig. 2(b)) results in a greater total extension $(\sum_{i=1}^{N-1}\Phi_{i})$ of the particle chain. Accordingly, the axial Coulomb forces exerted on the outermost particles (the first and the last particle of the particle chain) prevail for $N+1$ particles even at lower values of $\Phi$ compared to the $N$ particle chain. This ultimately leads to the flattening of the last potential well when $N$ is increased, as marked by the black crosses in Fig. 2(b). Nonetheless, this attribute is not entirely generic, as in the transition from $N=2$ to $N=3$, two minima disappear, and for $6\leq N\leq 8$, respectively $9\leq N\leq 14$, the last minimum is stable. By increasing the geometric parameter $\rho$ of the helix, the last potential well of each curve in Fig. 2(b) would be shifted to smaller values of $\Phi$ and the complex behavior of the intersection for $0<\Phi<1$ would vanish and become a single potential well. This indicates a similar behavior of the many-body potential for an increase in $\rho$ compared to Fig. 2(a). It is imperative to reiterate that Fig. 2(b) is merely an intersection and does not embody the high-dimensional potential landscape. Nevertheless, this intersection partially reflects the structural peculiarities of the potential landscape and its local equilibria, as we shall see in the subsequent sections.
@@ III Methodology and computational approach
The statistical and correlation analysis of the spatial and energetic properties of the ECs of a long-range interacting particle system on a helix necessitates the finding of all local minima of our high-dimensional oscillatory potential energy surface (see Eq. (4)). This is illustrated in Fig. 2(b) and will be further elaborated below: the potential energy landscape shows a mix of regular and irregular oscillations and undergoes substantial fluctuations across various particle configurations while maintaining a hierarchical structure. The exponential growth in configuration space volume, concomitant with the rise in dimensionality, poses a major challenge in identifying all local equilibria of the potential landscape. We address this by implementing a multi-start approach, entailing the definition of multiple sampling points. Each sampling point is then utilized as an initial condition for a local optimizer. Notwithstanding the aforementioned ”disorder” inherent in the potential landscape, the repulsion within and across helical windings must engender approximate regularity in the occurrence of ECs, which motivates the definition of the sampling points as a periodic hyperlattice $L\subseteq Q$ in the configuration space $Q\subseteq\mathbb{R}^{N-1}$. Every point in this space is given by a vector $\mathbf{\Phi}=(\Phi_{1},\Phi_{2},\ldots,\Phi_{N-1})$ of the internal coordinates. The hyperlattice is defined as $L=\tilde{\Phi}_{1}\times\tilde{\Phi}_{2}\times\ldots\times\tilde{\Phi}_{N-1}$, where the subsets $\tilde{\Phi}_{k}\subseteq\mathbb{R}$, with $k=1,2,\ldots,N-1$, define $n$ evenly spaced points in the $k$-th dimension in a given interval $a,b]\subseteq\mathbb{R}$. While $a=0.01$ was fixed, $b$ was chosen in the range $4\leq b\leq 10$, depending on $N$. The optimization algorithm employed is L-BFGS-B [39, 40], which implements a quasi-Newton method with bound constraints for the $k$-th coordinate.
@@
The computational effort resulting from this approach is increasing exponentially with the number of particles employed in the system ($\propto n^{N-1}$). Furthermore, the highly oscillatory structure of the potential landscape requires a high sampling point density $\delta=n/(b-a)$ to find all local equilibria. Reducing the search space is one way to lower the computing costs under these conditions. As indicated in the previous sections, the maximum value of $\Phi_{k}$ that allows for ECs either remains approximately constant or decreases when a particle is added to the chain. Consequently, this maximum value serves as an upper bound constraint $b$ for the system with an incrementally larger number of particles. On the other hand, the inversion of the helix with a given configuration $\mathbf{\Phi}$ results in an identical configuration $\mathbf{\Phi}$ with the same potential energy. Therefore, the search space can be further reduced by approximately half through the elimination of these identical configurations. We accomplish this by employing a recursive constraint for $\Phi_{i}$ and $\Phi_{N-i}$, $i=1,\ldots,k=\lfloor(N-1)/2\rfloor$ and $N\geq 3$. We enforce that $\Phi_{1}\geq\Phi_{N-1}$; $\Phi_{2}\geq\Phi_{N-2}$ if $\Phi_{1}=\Phi_{N-1}$; $\Phi_{3}\geq\Phi_{N-3}$ if $\Phi_{1}=\Phi_{N-1}\land\Phi_{2}=\Phi_{N-2}$; $\ldots$ ; $\Phi_{k}\geq\Phi_{N-k}$ if $\Phi_{1}=\Phi_{N-1}\land\,\ldots\,\land\Phi_{k-1}=\Phi_{N-(k-1)}$. Formally, this constraint can be written as a lexicographical ordering of the subsets $A=\left\{\Phi_{1},\ldots,\Phi_{k}\right\}$ and $B=\left\{\Phi_{N-1},\ldots,\Phi_{N-k}\right\}$: $A\geq_{\mathrm{lex}}B$. In our optimization algorithm, points of $L$ that invalidate this constraint are introduced into the interstitial voids of the lattice structure.
@@ IV.1 Spatial equilibrium configurations: Statistics and correlations
Figure 3: The total number of attained ECs, contingent upon the number of particles. The inset illustrates the number of ECs by $N$ on a double logarithmic scale, exhibiting that it follows a power law $(\approx N^{2.7})$.
The analysis of the statistics and correlations of the spatial properties of the ECs will be conducted for each particle number in the range 2-8. In consideration of the denotation of an EC as $\mathbf{\Phi}_{0}=(\Phi_{0}^{1},\Phi_{0}^{2},\ldots,\Phi_{0}^{N-1})$, the distribution of the totality of the interparticle distances $\Phi_{0}^{i}$, $i=1,2,\ldots,N-1$ of all ECs will be examined, independent of their sequential ordering within an EC. Building upon this distribution, a mapping of these $\Phi_{0}^{i}$ to distinct clusters will be employed to elucidate the spatial structure of the ECs. This will be achieved by indexing the clusters and then correlating the $\Phi_{0}^{i}$ values within an EC based on their assigned cluster index. Finally, we will gain insight into the spatial structure by studying the distribution of ECs within the (N1)(N-1)-dimensional potential landscape, quantifying their Euclidean distances in the configuration space.
@@ IV.1.1 Distribution of interparticle distances
Figure 5: (a) The probability on how many distinct clusters the interparticle distances $\Phi_{0}^{i}$ of an EC (Clusters/EC) are distributed in dependence of the number of particles $N$. For example, for the EC $\mathbf{\Phi}_{0}\approx(1.35,0.36)$, $C(\Phi_{0}^{1})=2$ and $C(\Phi_{0}^{2})=1$, so $\Phi_{0}^{i}$ is distributed on two Clusters in that EC. The probability is normalized for each NN. (b-h) The probability on which specific clusters the $\Phi_{0}^{i}$ are distributed, given the value of Clusters/EC. Note that $C(\Phi_{0}^{i})$ is cut at the value of 8 because $C(\Phi_{0}^{i})=9,10$ are only relevant for the trivial case of $N=2$. The probability is normalized for each value of Clusters/EC.
@@
A generally similar but more complex transition can be observed for the Clusters 1 and 2 (see Fig. 4(a-g)). While the maximum value of $\Phi_{0}^{i}$ in these clusters also initially increases, the minimum value of $\Phi_{0}^{i}$ is decreasing, thereby forming a fine structure with two subclusters. For Cluster 2 the upper subcluster is broader than the lower one, evidenced by an asymmetry in the corresponding probability distributions. For $N>5$, cluster 2 decreases, predominantly by the fading of the lower subcluster. In contrast, cluster 1 exhibits a less pronounced dominance of the upper subcluster, characterized by a more equitable distribution of $\Phi_{0}^{i}$ across both clusters. This results in probability distributions, that are more or less symmetrical. Additionally, cluster 1 merely shows an increase in its width, while a subsequent decrease may only commence at higher values of $N$.